OpenStudy (mrbee):

What is the solution to log25x =3?

OpenStudy (zzr0ck3r):

Is it \(\log(25x))\) or \(\log_{25}(x)\)?

jhonyy9 (jhonyy9):

what is the base of this logarithm Bee ?

jhonyy9 (jhonyy9):

or this mean log base 2 of 25x = 3 ?

OpenStudy (mrbee):

The answers can be x=3/25 | x=8/25 | x=35 | x=40

jhonyy9 (jhonyy9):

ok i understand but we need knowing the base of logarithm

OpenStudy (mrbee):

10

OpenStudy (mrbee):

40

jhonyy9 (jhonyy9):

so than is 10 how can you rewriting the right side using log base 10 ?

OpenStudy (mrbee):

Thanks I got it. x=40

jhonyy9 (jhonyy9):

are you sure ?

jhonyy9 (jhonyy9):

do you can proving it ?

jhonyy9 (jhonyy9):

log 25x =3 so this mean that 10^3 = 25x using this propertie of logarithms that log(a) b = x => a^x = b hope helped

jhonyy9 (jhonyy9):

Sila do you agree it ?

OpenStudy (zzr0ck3r):

correct

OpenStudy (sila1453):

yes